# PASS-FC: Progressive and Adaptive Search Scheme for Fact Checking of Comprehensive Claims

### v1

@@ Abstract
Automated fact-checking faces challenges in handling complex real-world claims. We present PASS-FC, a novel framework that addresses these issues through claim augmentation, adaptive question generation, and iterative verification. PASS-FC enhances atomic claims with temporal and entity context, employs advanced search techniques, and utilizes a reflection mechanism. We evaluate PASS-FC on six diverse datasets, demonstrating superior performance across general knowledge, scientific, real-world, and multilingual fact-checking tasks. Our framework often surpasses stronger baseline models. Hyperparameter analysis reveals optimal settings for evidence quantity and reflection label triggers, while ablation studies highlight the importance of claim augmentation and language-specific adaptations. PASS-FC’s performance underscores its effectiveness in improving fact-checking accuracy and adaptability across various domains. We will open-source our code and experimental results to facilitate further research in this area.
@@ 1 Introduction
The proliferation of online information and the advent of large language models (LLMs) have significantly increased the volume and complexity of content available to users (Tian et al., 2024; Wang et al., 2024a). This surge in information has brought the critical task of fact-checking to the forefront (Huang et al., 2022). Standard approaches for automated fact-checking comprise three stages (Min et al., 2023; Chern et al., 2023; Wei et al., 2024; Setty and Setty, 2024): (1) breaking down the content into atomic claims [^1] An atomic claim is a short sentence conveying a single piece of information, as defined by Factscore (Min et al., 2023).; (2) conducting web-based searches to gather relevant evidence; and (3) verifying each claim against the retrieved evidence. However, these methods struggle to effectively address certain types of issue, as illustrated in Figure 1.
Firstly, the conventional definition of atomic facts overlooks the potential ambiguities in the claim date and entity references within the context of world knowledge (Gunjal and Durrett, 2024; Chiang and Lee, 2024). As illustrated in the left path of Figure 1, the Madagascar Zone at Universal Studios Singapore was once operational but had closed at the time of verification. This temporal nuance is missed by the atomic claim. The absence of precise temporal and entity specifications can lead to validation against evidence that, while seemingly relevant, does not accurately reflect the current state of affairs.
@@
Figure 1: The workflow comparison between traditional fact-checking pipelines and PASS-FC, which enhances atomic claims by incorporating temporal and entity information, and utilizes advanced search and multilingual search to obtain relevant and sufficient evidence. The content within the rectangular boxes represents the retrieved evidence.
Secondly, the question generation module faces inherent challenges due to the uncertainty of retrieval outcomes (Lin et al., 2023). Search engines encompass vast and complex repositories of world knowledge, making it difficult to guarantee that a simple retrieval attempt based on existing information will yield relevant, credible, and sufficient evidence within the top results (Yu et al., 2024; Song et al., 2024). This ambiguity is particularly pronounced in information-seeking scenarios, where users formulate queries about topics they are unfamiliar with (Park et al., 2021). The complexity of real-world information often requires multiple retrieval iterations (Khattab et al., 2021; Zhang et al., 2024) and sophisticated query refinement strategies to bridge the gap between initial queries and the desired evidence (Ousidhoum et al., 2022; Schlichtkrull et al., 2023; Chen et al., 2022). As illustrated in Figure 1, these strategies include the use of advanced search operators and multilingual retrieval techniques, which fully leverage search engine capabilities to enhance the efficiency and accuracy of the verification process.
In this paper, we propose PASS-FC, a framework that iteratively performs adaptive question generation and verification on decontextualized atomic facts. Specifically, PASS-FC first enhances the atomic facts by supplementing them with occurrence times and entity descriptions to prevent subsequent ambiguities (3.1.2). In the question generation module, PASS-FC incorporates advanced search and multilingual search capabilities, allowing the model to adjust its retrieval strategies based on the claim and previous feedback (3.2). Finally, the model reflects on the entire process and determines whether to initiate another round of retrieval, verification, or to conclude the process (3.5). Upon completion, PASS-FC returns a factuality label, providing a final assessment of the claim’s veracity based on the comprehensive evidence gathered.
We conducted extensive testing on six diverse datasets, demonstrating that PASS-FC substantially improves fact verification performance, even outperforming models with stronger base capabilities in some scenarios. Our framework shows remarkable versatility, excelling in general knowledge, scientific, real-world, and multilingual fact-checking tasks (4.1). Furthermore, we performed a detailed analysis of various hyperparameters, revealing intriguing insights such as the optimal evidence count and the impact of reflection mechanisms. Our experiments uncovered that different language models benefit differently from iterative processes (4.2). Lastly, our ablation studies quantified the contribution of each module, demonstrating the crucial roles of claim augmentation, advanced search techniques, and language-specific adaptations in multilingual settings (4.3).
The main contributions of this paper are as follows: (1) We propose PASS-FC, a novel framework that enhances atomic facts with temporal and entity context, and employs iterative, adaptive question generation and verification for improved fact-checking accuracy. (2) We demonstrate PASS-FC’s effectiveness across diverse fact-checking scenarios, including general knowledge, scientific, real-world, and multilingual tasks, often outperforming strong baseline models. (3) We provide insights into optimal fact-checking processes through comprehensive hyperparameter analysis and ablation studies, revealing the influence of evidence quantity, reflection mechanisms, and language-specific adaptations.
@@ 2 Related Work
Atomic Claim and Decontextualization Effective fact verification requires a clear definition of facts (Ni et al., 2024), often necessitating the identification of atomic claims within longer texts. Numerous studies (Hu et al., 2024; Bayat et al., 2023; Min et al., 2023; Song et al., 2024) have proposed definitions for atomic facts. Chiang and Lee (2024) first highlighted the issue with atomic facts in which excessive atomization can lead to entity ambiguity. They addressed this problem through entity linking. Gunjal and Durrett (2024) introduced the concept of molecular facts, refining the approach by expanding entity and event descriptions within atomic facts. While these methods have shown promise, they have primarily been tested in biographical generation tasks. Our work distinguishes itself in several ways: Beyond supplementing entity descriptions, we also consider metadata such as temporal information that can contribute to ambiguity. Crucially, we extend beyond previous studies by evaluating the impact of these enhancements on fact verification in real-world scenarios.
@@
Question Generation for Fact Verification To address real-time information demands (Fierro et al., 2024; Kasai et al., 2022) in real-world scenarios, many fact verification frameworks (Gao et al., 2022; Chen et al., 2023a; Sun et al., 2024) generate questions to retrieve relevant knowledge from search engines like Google. Most existing works (Chern et al., 2023; Song et al., 2024; Wang et al., 2024a) rely on few-shot examples to prompt LLMs to ask direct questions or inquire about entity attributes in claims (Zhang et al., 2024; Wang and Shu, 2023). Schlichtkrull et al. (2023) demonstrates that human questioning strategies in fact verification are highly diverse, with an average similarity of only 0.25 between effective retrieval strategies across different individuals. Ousidhoum et al. (2022); Setty and Setty (2024) trained models to learn human-like questioning strategies based on existing datasets (Fan et al., 2020; Schlichtkrull et al., 2023; Chen et al., 2022; Park et al., 2021), but its domain is limited and lacks evaluation in real-world fact verification scenarios. Our work innovates by leveraging Google’s advanced search operators [^2] https://www.google.com/advanced_search combined with multilingual retrieval, aiming to obtain more useful evidence.
Iterative Verification Verifying complex claims often requires multiple iterations to reach accurate conclusions. SAFE (Wei et al., 2024) and KnowHalu (Zhang et al., 2024) explicitly allow models to perform multiple retrievals until they deem all useful information has been obtained. Other approaches Sun et al. (2024); Cohen et al. (2023) retrieve information only once but enable models to re-analyze verification results from different perspectives. ProgramFC (Pan et al., 2023) and Self-Checker (Li et al., 2024) treat all steps in the verification process as tools, allowing models to implicitly choose multiple retrievals or verifications by selecting appropriate tools at each step. Our work introduces a reflection step, which decides whether to retrieve more information, perform additional verification, or conclude the process after each verification cycle.
@@ 3 PASS-FC
Figure 2: Overview of the PASS-FC framework.
Figure 3: The timeline for the claim "Universal Studios features a Madagascar zone.". The valid period, shown as a green area, is marked with start and end dates from authorized sources. Three points (a, b, c) indicate claim dates inferred from the validation date (2024-12-21) and the claim’s time description: a. "2010" b. "three years ago" c. No time description. If an inferred date falls outside the valid period, the claim is unsupported for that time description.
We begin by formulating the fact-checking task. Given a query in any language, a response is generated either by a human or an LLM. The system’s objective is to determine the veracity of this response within the context of the query and any relevant metadata, such as the date.
PASS-FC addresses this task through a two-step process, as illustrated in Figure 2. In the first step, the system extracts n comprehensive claims $\{c_{1},c_{2},\ldots,c_{n}\}$ from the response. Subsequently, for each claim, PASS-FC employs an iterative process of query generation, evidence retrieval, and claim verification. At the conclusion of each iteration, PASS-FC evaluates the process thus far, determining whether to terminate or continue. If continuation is warranted, it generates appropriate tools and instructions for the subsequent round. Prior to initiating the fact-checking process, PASS-FC implements a Language-Specific Initialization step. This phase involves configuring the model to think and operate in the source language specified by the user. By default, the source language is set to English, but it adapts to other languages based on user input.
Notably, this process operates independently of gold evidence. Each phase of PASS-FC is implemented through carefully crafted prompts to an LLM. Detailed example prompts are available in Appendix A.6. For psedocode of PASS-FC, please refer to algorithm 1.
@@ 3.1 Claim Detection
Claim detection is a crucial preliminary step in our fact-checking process, involving claim decomposition and augmentation. This step facilitates fact-checking by breaking down the response into suitable decontextualized atomic facts.
@@ 3.1.1 Claim Decomposition
Long-form responses often contain multiple pieces of information, making it impractical to fact-check them as a whole. Splitting the response into atomic facts has been widely recognized as an effective approach (Min et al., 2023; Chern et al., 2023). Drawing from recent NLP research (Ni et al., 2024; Gunjal and Durrett, 2024; Min et al., 2023), we define an atomic fact as a statement or assertion that: (1) can be objectively verified as true or false based on empirical evidence or reality, (2) represents a single, indivisible unit of information, (3) is self-contained and context-independent
Given an input response r with its corresponding prompt p, we decompose the response into a set of claims $\{c_{1},c_{2},\ldots,c_{n}\}$ that adhere to this definition. To ensure the atomicity of claims, we employ an iterative process: if a claim exceeds a predefined word threshold k [^3] Empirically predefined length threshold. Set to 10 for English., it undergoes further decomposition into simpler sub-claims. This approach strikes a balance between efficiency and effectiveness, resulting in verifiable, atomic, and self-contained claims. Detailed prompts is shown in Figure 5 and Figure 6.
3.1.2 Claim Augmentation
Real-world claims often possess complex and nuanced contexts. Although we generate self-contained claims, they require decontextualization to avoid temporal and entity ambiguities in world knowledge. Claim augmentation addresses this by appending a claim period and entity brief to each atomic claim, forming a comprehensive claim as illustrated in Figure 2.
Claim Period
We introduce the claim period to enhance the temporal context of factual claims. It represents the time span during which the event or situation described in the claim occurred or was true. To determine the claim period, our method analyzes explicit and implicit time references within the claim text, as well as metadata provided alongside the claim. Figure 3 demonstrates how variations in the claim’s temporal description affect the corresponding claim period. A claim is considered true if its claim period falls within the time span covered by supporting evidence.
Entity Brief
The entity brief is a crucial component of claim augmentation that focuses on uniquely specifying each entity referenced in a claim (Fan et al., 2020). Using the given prompt, response, and metadata, a language model reasons to provide concise descriptions for every entity mentioned. Unlike approaches such as Molecular Facts (Gunjal and Durrett, 2024), which only describe potentially ambiguous entities like person or place names, our method includes all entities in the brief. This comprehensive approach acknowledges the vast and dynamic nature of world knowledge, recognizing that entity ambiguity may only become apparent during the evidence-searching process. By fixing entity references within the fact-checking context, the entity brief ensures consistency throughout the verification process.
Figure 2 provides an example of how a claim is augmented with claim period and entity briefs, which work in tandem to identify relevant evidence pertaining to the same entities and time frame mentioned in the claim.
With the atomic claim augmented with a claim period and entity brief, PASS-FC proceeds to iteratively perform query generation, evidence retrieval, claim verification, and reflection. The prompt, response, and metadata no longer serve as input context in the subsequent process. From this point forward, we refer to the augmented atomic claim as the comprehensive claim.
3.2 Query Generation
Our query generation module enhances the capability of LLMs to retrieve relevant, sufficient, and trustworthy evidence. This module processes a comprehensive claim along with any preceding step history to generate effective search queries. These queries are then used to extract evidence from search engines, facilitating the fact-verification process. The selection of specific search tools is determined by the reflection module based on previous history, with the advanced search tool being the default choice for the initial verification attempt.
3.2.1 Advanced Search
Advanced search operators are specialized commands and syntax employed by various search engines to refine and focus search results. These operators enable users to construct more precise queries by filtering information, specifying exact phrases, excluding certain terms, and combining multiple search criteria. Common examples include quotation marks for exact matches, Boolean operators (AND, OR), minus signs for exclusion, wildcards, and parentheses for grouping terms.
We incorporate an advanced search tool into our query generation module, leveraging these operators to create more targeted and effective search queries for fact verification. For each claim, our system generates two carefully crafted queries, utilizing appropriate operators to enhance the retrieval of relevant and reliable evidence. For a detailed explanation of the advanced search operators used in our system, please refer to Figure 8 in A.6, which contains the complete advanced search prompt.
3.2.2 Site-restricted Search
The site-restricted search tool is a crucial component of query generation module, designed to refine the scope of evidence retrieval. This tool enhances the fact-checking process in several ways. Firstly, it generates a list of credible domain suffixes for focused searching while identifying and excluding unreliable sources. This process leverages the LLM’s extensive knowledge of domain names and incorporates insights from previous search results. Secondly, the tool offers flexibility by accommodating user input. Users can manually specify preferred domains (e.g., wiki.org), ensuring that all fact-checking evidence is sourced from a curated set of trusted websites. This feature allows for customized, domain-specific searches that align with user preferences.
The tool’s output is seamlessly integrated with the advanced search queries, effectively narrowing the search to more reliable and relevant sources. Site-restricted search tool operates in parallel with the advanced search tool. While the advanced search tool can be used independently to generate queries without domain restrictions, the site-restricted search tool, when employed, works in conjunction with advanced search operators. This combination allows for more targeted queries that not only utilize advanced search techniques but also focus on specific, credible domains.
By integrating domain restrictions with advanced search operators, this approach directs searches to authoritative sources while maintaining the flexibility to perform broader searches when necessary. This targeted yet flexible method significantly enhances both the efficiency and accuracy of the fact-checking process, ensuring that the evidence gathered is not only relevant but also from credible sources.
3.2.3 Multilingual Search
To enhance the global applicability of PASS-FC, we introduce a multilingual search tool that expands the scope of evidence gathering beyond the user-defined source language. This tool analyzes the given claim to identify up to two relevant languages, in addition to the source language, from a predefined list of 46 languages supported by Google Search [^4] https://support.google.com/googleplay/android-developer/table/4419860?hl=en&sjid=11904773475773808427-AP. The source language, which defaults to English if not specified by the user, serves as the primary language for all steps in the fact-checking process.
The multilingual search tool considers key elements such as locations, people, news sources, and event places to determine which additional languages might offer more comprehensive, accurate, or up-to-date information. This approach is based on the principle that local language sources often provide more detailed and nuanced coverage of events or topics, particularly for claims involving international events, cultural phenomena, or region-specific topics.
Once these languages are identified, the system re-engages the advanced search query generation process, creating new queries in each of the selected languages. This crucial step ensures that the search terms are appropriately localized and culturally relevant, maximizing the effectiveness of the multilingual search.
3.3 Evidence Retrieval
Once queries are generated, they are submitted to a commercial search engine API [^5] https://serper.dev/ to collect relevant evidence. While multiple search engines support the advanced techniques we employ, we have chosen Google Search for its superior quality and comprehensive coverage. For each query, we retrieve the top-k (e.g., k=10) search results. From these results, we extract the title, snippet, and URL of each item. These elements are then combined to create a consolidated set of evidence for further analysis.
3.4 Claim Verification
The claim verification stage assesses whether a given claim is supported, contradicted, or inconclusive based on the available evidence. This process evaluates the consistency between the claim and the evidence, considering factual details, temporal aspects, and source credibility. The verification process is iterative, incorporating feedback from previous attempts to refine its judgment. We use three predefined veracity labels: supported (the majority of evidence corroborates the claim), contradicted (the majority of evidence opposes the claim), and inconclusive (either no relevant evidence is found, or there is conflicting evidence from credible sources that is temporally consistent with the claim).
3.5 Reflection
The reflection module analyzes the fact-checking process and determines whether further iteration is necessary. It provides textual feedback on the entire process and makes informed decisions based on historical steps and the current state.
When issues are identified, the module selects appropriate tools and generates feedback for improvement. If additional information is needed, it employs retrieval tools from the question generation module to refine search queries or explore specific credible domains. When key information has been overlooked, the module recommends revisiting the claim verification step, providing feedback that alters the model’s reasoning approach.
The process concludes when the historical steps are deemed satisfactory or after a predetermined number of iterations. This adaptive mechanism ensures continuous improvement in fact-checking accuracy while maintaining efficiency.
3.6 History Management
Our history management approach adapts to the context length capabilities of different language models. For models with context lengths of 8,000 tokens or more, we maintain a dictionary containing all historical information, including brief descriptions and results of each step. For models with limited context lengths, such as GPT-3.5-Turbo (4,096 tokens), we retain only the comprehensive claim, the previous iteration’s feedback, and the current round’s history. This strategy optimizes performance across various language models while preserving essential contextual information.

### v2

@@ Abstract
Automated fact-checking (AFC) still falters on claims that are time-sensitive, entity-ambiguous, or buried beneath noisy search-engine results. We present PASS-FC, a Progressive and Adaptive Search Scheme for Fact Checking. Each atomic claim is first grounded with a precise time span and disambiguated entity descriptors. An adaptive search loop then issues structured queries, filters domains through credible-source selection, and expands queries cross-lingually; when necessary, a lightweight reflection routine restarts the loop. Experiments on six benchmarks—covering general knowledge, scientific literature, real-world events, and ten languages—show that PASS-FC consistently outperforms prior systems, even those powered by larger backbone LLMs. On the multilingual X-FACT set, performance of different languages partially correlates with typological closeness to English, and forcing the model to reason in low-resource languages degrades accuracy. Ablations highlight the importance of temporal grounding and the adaptive search scheme, while detailed analysis shows that cross-lingual retrieval contributes genuinely new evidence. Code and full results will be released to facilitate further research.
[…]
@@ 1 Introduction
The Web and large language models (LLMs) have created an information deluge, but also a minefield of inaccuracies (Tian et al., 2024; Wang et al., 2024a). This surge in information has brought the critical task of fact-checking to the forefront (Huang et al., 2022a). Standard approaches for automated fact-checking comprise three stages (Chern et al., 2023; Wei et al., 2024; Setty and Setty, 2024): (1) decomposing text into atomic claims [^1] An atomic claim is a short sentence conveying a single piece of information, as defined by Factscore (Min et al., 2023).; (2) searching the Web for relevant evidence; and (3) verifying each claim against that evidence. Unfortunately, two persistent bottlenecks still limit their reliability (Figure 1).
@@
Figure 1: Workflow comparison between a traditional fact-checking pipeline (left) and PASS-FC (right). PASS-FC enriches atomic claims with temporal and entity information, then employs advanced and multilingual search to collect relevant and sufficient evidence. Text inside rectangles shows retrieved snippets.
Bottleneck1: missing context.
The canonical atomic claim omits when a statement is intended to hold and which real-world entity it references (Gunjal and Durrett, 2024; Chiang and Lee, 2024). The left half of Figure 1 illustrates the pitfall: "Universal Studios features a Madagascar Zone." was once true but false in 2024, yet a date-free checker may still retrieve outdated corroboration (Gunjal and Durrett, 2024; Chiang and Lee, 2024).
Bottleneck 2: brittle search.
Question-generation modules in AFC must operate under uncertain retrieval outcomes (Lin et al., 2023). Vast, heterogeneous Web indices make it unlikely that single, naïve queries will surface relevant, credible, and sufficient evidence (Yu et al., 2024; Song et al., 2024). Information-seeking scenarios often require multiple iterations (Khattab et al., 2021; Zhang et al., 2024), advanced operators, domain filters, or multilingual queries before trustworthy sources emerge (Chen et al., 2022; Ousidhoum et al., 2022; Schlichtkrull et al., 2023) (as in the right path of Figure 1). Existing systems (Wei et al., 2024; Li et al., 2024) typically rely on simple query generation and hope that subsequent reflection alone can instruct it to uncover a workable verification path.
We address these challenges with PASS-FC. Each atomic fact is augmented with a temporal span and unique entity descriptors (§3.1.2). Our adaptive search scheme3.2) then crafts structured queries, selects credible domains, and, when needed, expands the search into other languages, progressively shrinking the evidence gap. After every round a reflection module (§3.3) inspects the verdict and either stops or issues a new tool plan. The system finally outputs a veracity label plus a concise rationale.
We first tune hyperparameters on a separate development set, revealing insights such as the optimal number of evidence passages and the varying benefits of reflection for different base models (Figure 4c). With these settings, PASS-FC exceeds all baselines on six benchmarks, even outperforming systems that use stronger backbone LLMs (§4.1). On the multilingual X-FACT set, performance of different languages partially correlates with typological distance from English, and forcing the model to reason in low-resource languages reduces accuracy (Table 7). Ablations quantify the impact of each module, underscoring the importance of temporal grounding and the adaptive search scheme (§4.2). Further analysis shows that foreign-language evidence adds genuinely new support (§4.3 and §4.4).
Contributions
  • We propose PASS-FC, the first framework that unifies temporal/entity grounding with an adaptive, multilingual search loop for fact-checking.
  • Extensive experiments demonstrate state-of-the-art accuracy on general, scientific, real-world, and multilingual benchmarks, even when PASS-FC uses smaller backbone LLMs.
  • We release code, prompts, and complete results, and provide practical insights into evidence quantity, reflection triggers, multilingual fact-checking, and the role of cross-lingual evidence.
@@ 2 Related Work
Models Structured Query Generation Credible-Source Selection Cross-Lingual Expansion Temporal Grounding Reflection
FacTool (Chern et al., 2023) - - - - -
KnowHalu (Zhang et al., 2024) - - - - $\checkmark$
FOLK (Wang and Shu, 2023) - search in wikipedia - - -
FactScore (Min et al., 2023) - - - - -
ProgramFC (Pan et al., 2023) - - - - $\checkmark$
Self-Checker (Li et al., 2024) - - - - $\checkmark$
SAFE (Wei et al., 2024) - - - - $\checkmark$
VERISCORE (Song et al., 2024) - - - - $\checkmark$
Hiss (Zhang and Gao, 2023) - - - - $\checkmark$
PACAR (Zhao et al., 2024) - - - - $\checkmark$
Rafts (Yue et al., 2024) - - - - -
MiniCheck (Tang et al., 2024) - - - - -
FactCheckGPT (Wang et al., 2024a) - - - - $\checkmark$
DEFAME (Braun et al., 2024) - - - - $\checkmark$
FIRE (Xie et al., 2024) - - - - $\checkmark$
PASS-FC (Our Method) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Comparison of fact-check frameworks across various capabilities. We claim all of them except reflection as our initial contribution to the automatic fact-checking (AFC) task.
@@
Atomic Claim and its Grounding Effective fact verification requires a clear definition of facts (Ni et al., 2024), often necessitating the identification of atomic claims within longer texts. Numerous studies (Hu et al., 2024; Bayat et al., 2023; Min et al., 2023; Song et al., 2024) have proposed definitions for atomic facts. Chiang and Lee (2024) first highlighted the issue with atomic facts in which excessive atomization can lead to entity ambiguity. They mitigate this with entity linking. Gunjal and Durrett (2024) introduced the concept of molecular facts, refining the approach by expanding entity and event descriptions within atomic facts. While these methods have shown promise, they have primarily been tested in biographical generation tasks. Our work distinguishes itself in two ways: (1) Beyond supplementing entity groundings, we also consider temporal grounding: an explicit date or time span that anchors truth to the correct period, preventing evidence drift. (2) We extend beyond previous studies by evaluating the impact of these enhancements on end-to-end fact verification in real-world scenarios. Table 1 shows that PASS-FC is the first system that brings temporal grounding into the AFC pipeline.
@@
Query Generation for Fact Verification To address real-world information demands (Fierro et al., 2024; Kasai et al., 2022), most frameworks (Gao et al., 2022; Chen et al., 2023a; Sun et al., 2024) generate queries to retrieve relevant knowledge from search engines like Google. Typical methods (Chern et al., 2023; Song et al., 2024; Wang et al., 2024a) rely on few-shot examples to prompt LLMs to ask direct or entity attributes-based questions (Zhang et al., 2024; Wang and Shu, 2023). Yet human studies (Schlichtkrull et al., 2023) reveal large variance in good queries, with an average similarity of 0.25 between effective retrieval strategies across different individuals. Learning that diversity helps (Ousidhoum et al., 2022; Setty and Setty, 2024), but existing datasets (Fan et al., 2020; Schlichtkrull et al., 2023; Chen et al., 2022; Park et al., 2021) cover only narrow domains. PASS-FC enforces three constraints before any query is sent: (i) Credible-Source Selection keeps only trustworthy domains, (ii) Cross-Lingual Expansion adds languages likely to contain fresh evidence, and (iii) Structured Query Generation composes Boolean queries that bind the claim, time, sources, and languages together. This adaptive search scheme is unique among current AFC systems (Table 1).
@@ 3 PASS-FC
Figure 2: Overview of the PASS-FC framework. We claim all of the procedures highlighted in bold orange as our initial contributions to the automatic fact-check (AFC) task.
(a) I. Timeline of the Madagascar zone. The green bar shows its verified open period obtained from authoritative sources; any claim period outside this bar is unsupported.
Claim Claim Period
a. Universal Studios features a Madagascar zone in 2010. 2010
b. Universal Studios features a Madagascar zone three years ago. 2021-12-21
c. Universal Studios features a Madagascar zone. 2024-12-21
(b) II. Mapping different temporal formulations (a, b, c) to claim periods given the validation date 2024-12-21.
Figure 3: Temporal grounding example for the claim “Universal Studios features a Madagascar zone.”
Task formulation.
Given a user prompt $p$ (in any language) and a free-form answer $r$ produced by either a human or an LLM, the goal is to decide whether the information in $r$ is true, false, or unverifiable. All judgments are made with respect to $p$ and the accompanying metadata (e.g. validation date).
PASS-FC workflow.
Figure 2 shows the pipeline. (1) We first split $r$ into a set of comprehensive claims $\{c_{1},\ldots,c_{n}\}$. (2) Each claim then enters an iterative loop composed of adaptive search, verification, and reflection (§3.2). After every round reflection either stops the loop or issues new tool instructions for the next cycle.
@@ 3.1 Claim Detection
This front-end stage converts a long answer into self-contained atomic facts and augments them with essential context.
@@ 3.1.1 Claim Decomposition
Following Ni et al. (2024); Gunjal and Durrett (2024); Min et al. (2023), an atomic fact is a minimal statement that (i) is objectively checkable, (ii) conveys exactly one proposition, and (iii) remains intelligible without extra context. Given $(p,r)$, an LLM decomposes $r$ into $\{c_{1},\ldots,c_{n}\}$ that satisfy the above criteria (prompt in Fig. 8).
3.1.2 Contextual grounding.
Even a self-contained claim may hide ambiguities about when it holds and which real-world entities it mentions. Therefore we append two concise fields to every atomic fact:
Entity Grounding
A unique cue phrase or descriptor for each entity referenced. It focuses on uniquely specifying each entity referenced in a claim (Fan et al., 2020). Using the given prompt, response, and metadata, a language model reasons to provide concise descriptions for every entity mentioned.
Temporal Grounding
The goal of temporal grounding is to attach a precise time span—called the claim period—to every atomic claim. The claim period is the interval in which the statement must hold for the claim to be true. Our procedure is purposely simple and deterministic, yet robust to the mismatch between the dataset’s validation date and the date on which the LLM is queried.
  1. 1.
    Parse temporal cues (see Fig. 3 for an example):
    1. a)
      Absolute description — if the claim contains an explicit time span (e.g., “in 2010”), we take that span verbatim as the claim period.
    2. b)
      Relative description — if the claim uses a relative phrase (e.g., “three years ago”), we resolve it with respect to the validation date $t_{\text{val}}$ supplied by the dataset.
    3. c)
      No description — if no temporal cue appears, we default to $t_{\text{val}}$ itself.
  2. 2.
    Align the LLM’s notion of “today”:
    LLMs are aware of the real calendar date $t_{\text{now}}$, whereas claims were annotated on $t_{\text{val}}$. Querying the model directly at $t_{\text{now}}$ may introduce fact drift. We therefore rewrite $t_{\text{val}}$ as the literal token Now in the prompt, let the model reason under this neutral anchor, and—after generation—replace Now with the real validation date. This single post-processing step decouples inference from wall-clock time and empirically stabilises accuracy.
A claim is deemed supported only if its derived claim period is fully covered by the time span of the retrieved evidence. The procedure grounds every claim to a fixed temporal frame, removes ambiguity for both absolute and relative descriptions, and, importantly, decouples LLM inference from the moving target of the current date—a key step toward robust, time-aware fact checking. The prompt is showed at Figure 9
[…]
3.2 Adaptive Search Scheme
The Adaptive Search Scheme equips PASS-FC with a focused, progressive retrieval mechanism. Given a comprehensive claim and the full reasoning trace, it produces search queries that maximize both relevance and trustworthiness of the evidence. At each iteration the reflection module selects the most suitable search tool with the Structured Query Generation (SQG) tool being the default choice for the initial verification attempt.
3.2.1 Credible-Source Selection
Credible-Source Selection (CSS) restricts the search space to domains that are likely to host reliable information. Conditioned on the current claim and earlier results, the LLM proposes a concise allow-list and discards sites of dubious quality. For instance, Figure 2 shows that visitingsingapore.com—an official portal inferred from previous hits—is retained, while reddit.com is dropped because its content is user-generated. In contrast to prior work (Schlichtkrull et al., 2023; Wang and Shu, 2023) that hard-codes a site list, CSS is fully automatic yet remains user-controllable: a user may also inject preferred domains (e.g., *.wikipedia.org). The generated list is forwarded to SQG through the prompt in Figure 10.
3.2.2 Cross-Lingual Expansion
Cross-Lingual Expansion (XLE) widens the evidence net beyond the user-specified source language (English is only the default when the user gives no preference). Given the current claim and its metadata, XLE selects at most two extra languages from the 46 that Google Search supports [^2] https://support.google.com/googleplay/android-developer/table/4419860?hl=en&sjid=11904773475773808427-AP. Selection follows three adaptive rules:
Content. If the claim discusses a country or culture, XLE adds that locale’s dominant language (e.g., “Singapore”, Malay or Mandarin).
Source. If the claim’s provenance hints at another linguistic context—through the URL domain, speaker nationality, publication venue, or any metadata field—XLE adds that language even when the claim itself is written in the source language or its content does not trigger the former rule.
Diversity. When the first two rules still leave the evidence sparse, XLE adds one high-resource language that is highly indexed for the topic (e.g., Spanish, French). If earlier retrieval has already yielded sufficient evidence, this step is skipped.
The source language is always preserved. For each language, XLE calls SQG to generate new language-specific queries (prompt in Figure 11).
3.2.3 Structured Query Generation
SQG converts the selected domains, languages, and temporal/entity grounds into executable search strings. It exploits advanced operators—site:, quotation marks, Boolean connectors, exclusion symbols, wildcards, and parentheses—to sharpen or broaden the search as needed. These operators are specialized syntax employed by various search engines. For every claim, SQG emits two complementary queries: one precision-oriented and one recall-oriented. A complete description of the operator set and the prompt template appears in Figure 12.
3.3 Verification, Reflection, and History Management.
For every structured query, we call the Google Search API [^3] https://serper.dev/ and keep the top-$k$ results ($k{=}10$, title–snippet–URL). The verifier LLM inspects the claim plus this evidence set and returns one of three labels: supported (most credible, temporally aligned evidence agrees with the claim); contradicted (credible, temporally aligned evidence refutes it); inconclusive (relevant evidence is missing or mutually conflicting). After each round, a reflection routine decides whether to (i) stop, (ii) launch an additional search via CSS/XLE/SQG, or (iii) re-evaluate the same evidence when the reasoning appears wrong. A light-weight history manager stores the claim, current evidence, verdict, and feedback. With long-context models ($>$8k tokens) the full log is preserved; with shorter-context models (e.g., GPT-3.5-Turbo) we keep only the claim, the most recent iteration, and a concise summary of earlier steps. The loop terminates when reflection signals “stop” or when a preset iteration budget is exhausted.